OpenStudy (anonymous):

Does this table represent a linear function? X-values Y-values 3 5 4 9 5 17 6 29

OpenStudy (anonymous):

no

OpenStudy (amistre64):

as long as the difference in ys are the same

OpenStudy (anonymous):

to be a linear function y = mx + b

OpenStudy (amistre64):

17-9 = 8 29 - 17 = 12 id have to go with no

OpenStudy (anonymous):

From 3 to 4, x changes 1 and y changes 4. From 4 to 5, x changes 1 and y changes 8. For a linear function, if the change in x is the same, the change in y should also be the same.